# Planar site percolation via tree embeddings

### v5

@@ 3. Characterization of Critical Percolation Probability
The following technical lemmas (Lemmas 3.1 and 3.2) were proved in [31], based on an adaptation of arguments in [13].
@@
Proof of Corollary 1.9Let $\mathcal{A}_{f}$ be the event that the number of infinite 1-clusters is finite and nonzero. Let $p>p_{c}^{site}(G)$. It suffices to show that if $\mathbb{P}_{p}(\mathcal{A}_{f})>0$, then $p\geq p_{conn}$.
@@ 4. Embedded Forests
Suppose Setup 4.4 holds. Then
@@
Suppose Setup 4.4 holds. Let $x:=\zeta_{i_{1}}$ and $y:=\xi_{j_{1}}$. Assume without loss of generality that $x$ precedes $y$ along $l_{uw}$. Then the edge along $\partial_{L}R_{i_{1}}$ immediately preceding $\mathrm{first}(\partial_{L}R_{i_{1}},l_{uw}(x,y])$ cannot be on the right of $l_{uw}(x,y]$, and the edge along $\partial_{R}U_{j_{1}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{1}},l_{uw}[x,y))$ cannot be on the right of $l_{uw}[x,y))$. (The opposite order $y\prec x$ follows by reversing the orientation of $l_{uw}$ and exchanging left/right.)
@@
This is proved by the same Gauss–Bonnet cycle argument as in Lemma 4.6: choose the first/last forbidden boundary–geodesic intersection (with respect to the orientation of $l_{uw}$), build a simple cycle using the corresponding geodesic subsegment of $l_{uw}$ and the relevant boundary ray, identify the unique “key” face along that geodesic subsegment using the extremality built into the choices of $i_{1},j_{1},i_{k},j_{r}$ in Setup 4.4, and then apply (2.2) to obtain a contradiction. We omit the repetitive details. ∎
@@
Suppose Setup 4.4 holds. Then
@@
(C) follows from Lemmas 4.7 and 4.9. See Figure 4.2.
@@
and $p_{\mathrm{conn}}(\cdot)$ is defined in Definition 1.7.
@@
When $p\in\bigl({p_{c}^{site}}(G),\,1-p_{c}(T)\bigr)$, the existence of infinitely many infinite $1$-clusters follows from Lemmas 1.9 and 4.10.
@@ 5. From infinitely many ends to infinitely many infinite clusters
$\mathbb{P}_{p}$-a.s. there are infinitely many infinite $1$-clusters. Note that by Lemmas 1.9 and 4.10, when $p\in(\tfrac{1}{2},1-p_{c}^{site}(G_{*}))$ there are $\mathbb{P}_{p}$-a.s. infinitely many infinite 0-*-clusters.
@@
Let $G$ satisfy the assumptions of Definition 4.1 and assume Assumption 5.3 holds with some $\delta>0$. Then for every $\in\bigl(1-p_{c}^{site}(G_{*})-\delta,\ 1-p_{c}^{site}(G_{*})\bigr), p\mathbb{P}_{p}$-a.s. every infinite 0-*-cluster has infinitely many ends and no isolated ends. (This indeed implies that every infinite 0-*-cluster has $2^{\aleph_{0}}$ ends.)
@@
Assume for contradiction that with strictly positive probability,
(5.2) $Infinite\ 0-*-clusters\ have\ an\ isolated\ end.$
Then by finite energy with strictly positive probability there exists an $1$-ended infinite 0-*-cluster. For $v\in V$, let $\mathcal{E}_{v}$ be the event that $v$ belongs to a $1$-ended infinite 0-*-cluster. Then $\mathbb{P}_{p}(\bigcup_{v\in V}\mathcal{E}_{v})>0$, so by countability there exists $v_{0}\in V$ such that
@@
Step 4: Sprinkling and contradiction. Assume (LABEL:eq:l03-contr) holds with positive probability. On this event, only finitely many infinite 0-*-clusters have at least two ends, and all infinite 0-*-clusters have finitely many ends. For each multi-ended infinite 0-*-cluster, choose a finite vertex set whose removal separates its ends; opening the union of these finite sets (a finite modification) produces, by finite energy, with positive conditional probability a configuration in which every infinite 0-*-cluster is $1$-ended. Denote this event by $A$, so $\mathbb{P}_{p}(A)>0$.
Choose $p^{\prime}>p$ so that $q^{\prime}:=1-p^{\prime}>p_{c}^{site}(G_{*})$, and couple $\omega_{p},\omega_{p^{\prime}}$ by i.i.d. uniforms $(U_{v})_{v\in V}$. Conditional on $\omega_{p}$, passing from $p$ to $p^{\prime}$ performs independent thinning of the 0-vertices with retention probability $q^{\prime}/q<1$. On $A$, every infinite 0-*-cluster $\xi$ satisfies $p_{c}^{site}(\xi)=1$ by Step 3, hence after thinning no infinite 0-*-cluster survives almost surely. Therefore $\mathbb{P}_{p^{\prime}}(\text{no infinite $0$-$*$-cluster})\geq\mathbb{P}_{p}(A)>0$.
This contradicts $q^{\prime}>p_{c}^{site}(G_{*})$, since by the definition of $p_{c}^{site}$ the Bernoulli($q^{\prime}$) percolation on $G_{*}$ has an infinite 11-*-cluster almost surely (equivalently, $\omega_{p^{\prime}}$ has an infinite 0-*-cluster almost surely). Hence (5.2) has probability 0. $\mathbb{P}_{p}$-a.s. every infinite 0-*-cluster has $2^{\aleph_{0}}$ ends follows from the fact that no inifinite 0-cluster has an isolated end, and the argument on Page 76 of [5] (which is stated for quasi-transitive graphs, but works for locally finite graphs as well). ∎
@@
Fix $p$ in the stated interval. By Lemma 5.4, $\mathbb{P}_{p}$-a.s. every infinite 0-*-cluster has infinitely many ends.
Case 1: $G$ has finitely many ends. Choose a finite subgraph $K$ such that $G\setminus K$ has finitely many infinite components and each infinite component is $1$-ended. If every infinite 0-*-cluster has infinitely many ends, then for some $1$-ended infinite component $\mathcal{C}$ of $G\setminus K$, the intersection $\xi\cap\mathcal{C}$ contains a connected component with infinitely many ends in $\mathcal{C}$. By planarity, these ends must be separated by infinite $1$-clusters, hence there are infinitely many infinite $1$-clusters.
@@
Assume for contradiction that $\omega$ has only finitely many infinite $1$-clusters. Let $\Theta$ be the union of all infinite 0-*-clusters together with all *-edges joining two vertices in the same infinite 0-*-cluster. By Lemma 5.4, $\Theta$ has infinitely many ends in the plane almost surely, and hence ${\mathbb{R}}^{2}\setminus\Theta$ has infinitely many unbounded connected components.
@@ 6. Non-self-touching polygons
A linear isoperimetric estimate for triangulations with minimum degree at least $d$ (see, e.g., the argument of [38, Theorem 6]) yields
@@
Combining Lemmas 6.2 and 6.4, we may choose $\delta>0$ such that $p_{c}^{site}(G_{*})+\delta<1/\gamma_{G_{*}}(e)$ for all $e\in E_{*}$. In particular, Assumption 5.3 holds for graphs whose every finite face is a triangle and whose minimum degree is at least $7$.
@@ 7. Uniform Percolation with respect to Binary Trees
The argument adapts the “finite-energy” multiscale proof of [41, Lemma 1.1], replacing balls by embedded binary trees (thus avoiding bounded-degree assumptions); see also the proof of [21, Theorem 2.3].
@@
This is proved by the same duplication/exploration argument as [41, Theorem 1.1], replacing “balls” by “depth-$N$ embedded binary trees” and invoking Lemma 7.1 instead of ball-uniformity. The bounded-degree assumption in [41] is only used to work with balls; the binary-tree version avoids this. We omit the repeated details. ∎

### v6

@@ 4. Characterization of Critical Percolation Probability
The following technical lemmas (Lemmas 4.1 and 4.2) were proved in [19], based on an adaptation of arguments in [9].
@@
Proof of Corollary 1.4. Let $\mathcal{A}_{f}$ be the event that the number of infinite 1-clusters is finite and nonzero. Let $p>p_{c}^{site}(G)$. It suffices to show that if $\mathbb{P}_{p}(\mathcal{A}_{f})>0$, then $p\geq p_{conn}$.
@@ 5. Embedded Forests
Suppose Setup 5.4 holds. Then
@@
Suppose Setup 5.4 holds. Let $x:=\zeta_{i_{1}}$ and $y:=\xi_{j_{1}}$. Assume without loss of generality that $x$ precedes $y$ along $l_{uw}$. Then the edge along $\partial_{L}R_{i_{1}}$ immediately preceding $\mathrm{first}(\partial_{L}R_{i_{1}},l_{uw}(x,y])$ cannot be on the right of $l_{uw}(x,y]$, and the edge along $\partial_{R}U_{j_{1}}$ immediately preceding $\mathrm{first}(\partial_{R}U_{j_{1}},l_{uw}[x,y))$ cannot be on the right of $l_{uw}[x,y))$. (The opposite order $y\prec x$ follows by reversing the orientation of $l_{uw}$ and exchanging left/right.)
@@
This is proved by the same Gauss–Bonnet cycle argument as in Lemma 5.6: choose the first/last forbidden boundary–geodesic intersection (with respect to the orientation of $l_{uw}$), build a simple cycle using the corresponding geodesic subsegment of $l_{uw}$ and the relevant boundary ray, identify the unique “key” face along that geodesic subsegment using the extremality built into the choices of $i_{1},j_{1},i_{k},j_{r}$ in Setup 5.4, and then apply (3.2) to obtain a contradiction. We omit the repetitive details. ∎
@@
Suppose Setup 5.4 holds. Then
@@
(C) follows from Lemmas 5.7 and 5.9. See Figure 5.2.
@@
and $p_{\mathrm{conn}}(\cdot)$ is defined in (1.3).
@@
When $p\in\bigl({p_{c}^{site}}(G),\,1-p_{c}(T)\bigr)$, the existence of infinitely many infinite $1$-clusters follows from Lemmas 1.4 and 5.10.
@@ 6. From infinitely many ends to infinitely many infinite clusters
$\mathbb{P}_{p}$-a.s. there are infinitely many infinite $1$-clusters. Note that by Lemmas 1.4 and 5.10, when $p\in(\tfrac{1}{2},1-p_{c}^{site}(G_{*}))$ there are $\mathbb{P}_{p}$-a.s. infinitely many infinite 0-*-clusters.
@@
Let $G$ satisfy the assumptions of Definition 5.1 and assume Assumption 6.3 holds with some $\delta>0$. Then for every $\in\bigl(1-p_{c}^{site}(G_{*})-\delta,\ 1-p_{c}^{site}(G_{*})\bigr), p\mathbb{P}_{p}$-a.s. not every end of infinite 0-*-clusters is isolated. In particular, this implies that there exists an infinite 0-*-cluster with infinitely many ends.
@@
First of all, by the same arguments as on page 76 of [3], not every end of infinite 0-*-clusters is isolated implies that there exists an infinite 0-*-cluster with $2^{\aleph_{0}}$ ends; here $\aleph_{0}$ is the cardinality of any countably infinite set. Hence it suffices to show that not every end of infinite 0-*-clusters is isolated. Since this is a tail-event, the Kolmogorov 0-1 law implies that it has probability of either 0 or 1.
@@
Assume for contradiction that a.s.,
(6.2) every end of infinite 0-*-clusters is isolated.
Let $\overline{G}=(\overline{V},\overline{E})$ be a graph constructed from $G$ as follows: $\overline{V}$ is obtained from $V$ by adding an extra vertex in each finite face, $\overline{E}$ is obtained from $E$ by adding an extra edge between each vertex in $\overline{V}\setminus{V}$ (i.e. an added vertex in a finite face $F$) and a vertex on the boundary of the face $F$. Then $\overline{G}$ is an infinite, connected, locally finite graph that can be properly embedded into ${\mathbb{R}}^{2}$. For site configuration $\omega\in\{0,1\}^{V}$, define a site configuration $\overline{\omega}\in\{0,1\}^{\overline{V}}$ by letting $\overline{\omega}(v)=\omega(v)$ if $v\in V$ and $\omega(v)=0$ if $v\in\overline{V}\setminus V$. Let $\overline{G}_{\omega}$ be the induced subgraph of $\overline{G}$ by state-0 vertices in $\overline{\omega}$. Then infinite 0-*-clusters in $\omega$ are in 1-1 correspondence with infinite components in $\overline{G}_{\omega}$. When (6.2) holds for $\omega$, every end of $\overline{G}_{\omega}$ is isolated.
Since $\overline{G}_{\omega}$ is a locally finite planar graph, it admits a Freudenthal embedding into the sphere $\mathbb{S}^{2}$ (see Section 2 of [20]; see also Section 8.6 of [8], Proposition 1.22 in [23]); and one can define a topology on the union $|\overline{G}_{\omega}|$ of $\overline{G}_{\omega}$ and its ends, such that the Fredenthal embedding of $|\overline{G}_{\omega}|$ into $\mathbb{S}^{2}$ is a homeomorphism from $|\overline{G}_{\omega}|$ onto its image. Under this embedding, each end of $|\overline{G}_{\omega}|$ corresponds an accumulation point of the embedding. Since $\mathbb{S}^{2}$ is first countable and each end of $|\overline{G}_{\omega}|$ is isolated, in $\omega$ infinite 0-*-clusters have countably many ends if (6.2) holds.
Moreover, on (6.2) by finite energy, with strictly positive probability there exists at least one $1$-ended infinite 0-*-cluster. For $v\in V$, let $\mathcal{E}_{v}$ be the event that $v$ belongs to a $1$-ended infinite 0-*-cluster. Then $\mathbb{P}_{p}(\bigcup_{v\in V}\mathcal{E}_{v})>0$, so by countability there exists $v_{0}\in V$ such that
@@
Step 4: Sprinkling and contradiction. Assume (6.2) holds ${\mathbb{P}}_{p}$ a.s.
Choose $p^{\prime}>p$ so that $q^{\prime}:=1-p^{\prime}>p_{c}^{site}(G_{*})$, and couple $\omega_{p},\omega_{p^{\prime}}$ by i.i.d. uniforms $(U_{v})_{v\in V}$. Conditional on $\omega_{p}$, passing from $p$ to $p^{\prime}$ performs independent thinning of the 0-vertices with retention probability $q^{\prime}/q<1$. If ${\mathbb{P}}_{p^{\prime}}$-a.s. there exists an infinite 0-*-cluster, then with strictly positive probability there exists a finite set of vertices $K$ such that removing $K$, there is an 1-ended infinite 0-*-cluster at level $p$ including an infinite 1-*-cluster at level $p^{\prime}$. But this has probability 0 by Step 3 and the fact that the total number of ends (all isolated) of infinite 0-*-clusters is countable.
This contradicts $q^{\prime}>p_{c}^{site}(G_{*})$. More precisely, by the definition of $p_{c}^{site}$ the Bernoulli($q^{\prime}$) percolation on $G_{*}$ has an infinite 11-*-cluster almost surely (equivalently, $\omega_{p^{\prime}}$ has an infinite 0-*-cluster almost surely). Hence (6.2) has probability 0. ∎
@@
Fix $p$ in the stated interval.
Case 1: $G$ has finitely many ends. Choose a finite subgraph $K$ such that $G\setminus K$ has finitely many infinite components $H_{1},\ldots,H_{n}$ and each infinite component is $1$-ended. Then
$p_{c}^{site}(G_{*})=\min_{1\leq i\leq n}\{p_{c}^{site}([H_{i}]_{*})\}$
Let $H\in\{H_{1},\ldots,H_{n}\}$, such that $p_{c}^{site}(G_{*})=p_{c}^{site}(H_{*})$. Then $H$ is an infinite, connected, one-ended graph properly embedded into ${\mathbb{R}}^{2}$. We apply Lemma 6.4 to $H$, since $H$ is one-ended, ${\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in $H$. Since $H$ is a component obtained by removing a finite set of vertices $K$ from $G$, we infer that ${\mathbb{P}}_{p}$-a.s. there are infinitely many infinite 1-clusters in $G$.
@@
Assume for contradiction that $\omega$ has only finitely many infinite $1$-clusters. Let $\Theta$ be the union of all infinite 0-*-clusters together with all *-edges joining two vertices in the same infinite 0-*-cluster. By Lemma 6.4, ${\mathbb{R}}^{2}\setminus\Theta$ has infinitely many unbounded connected components.
@@ 7. Non-self-touching polygons
A linear isoperimetric estimate for triangulations with minimum degree at least $d$ (see, e.g., the argument of [24, Theorem 6]) yields
@@
Combining Lemmas 7.2 and 7.4, we may choose $\delta>0$ such that $p_{c}^{site}(G_{*})+\delta<1/\gamma_{G_{*}}(e)$ for all $e\in E_{*}$. In particular, Assumption 6.3 holds for graphs whose every finite face is a triangle and whose minimum degree is at least $7$.
@@ 8. Uniform Percolation with respect to Binary Trees
The argument adapts the “finite-energy” multiscale proof of [27, Lemma 1.1], replacing balls by embedded binary trees (thus avoiding bounded-degree assumptions); see also the proof of [14, Theorem 2.3].
@@
This is proved by the same duplication/exploration argument as [27, Theorem 1.1], replacing “balls” by “depth-$N$ embedded binary trees” and invoking Lemma 8.1 instead of ball-uniformity. The bounded-degree assumption in [27] is only used to work with balls; the binary-tree version avoids this. We omit the repeated details. ∎